A long thin solenoid has radius 2.40 cm and 1000 turns per meter. The current in the solenoid is increasing at a constant rate of 50.0 A/s. Determine the magnitude of the induced electric field at a point inside the solenoid and 0.80 cm from the axis of the solenoid.


1
Expert's answer
2022-01-18T15:13:08-0500

"B=\mu_0nI,"

"\frac{dB}{dt}=\mu_0n\frac{dI}{dt}=0.063~\frac Ts,"

"E=\frac l2\cdot\frac{dB}{dt}=2.5\cdot 10^{-4}~\frac Vm ."